# Optimizing quantum violation for multipartite facet Bell inequalities

### v1

@@ 22 $\sqrt{2}$ 2-2 (0,0,1,1,1)(0,0,1,1,-1)
We now apply our method to the $(N,3,2)$ scenario. Introducing an additional measurement enables the construction of Bell inequalities with larger ratios, i.e., $\Delta_{N,3}^{\max}\geq\Delta_{N,2}^{\max}$. Thus, these inequalities are more robust and may be easier to certify experimentally. For $m=3$, all tight Bell inequalities can be enumerated only when $N<5$ [46], and we list these inequalities in [44]. For larger $N$, a complete enumeration is infeasible, but our gradient-based method remains applicable for searching facet Bell inequalities. For small $N$ we find that $\Delta_{N,3}^{\max}=\Delta_{N,2}^{\max}$, while for $N\geq 4$ the scenario with three measurements indeed leads to a larger ratio $\Delta_{N,3}^{\max}\geq\Delta_{N,2}^{\max}$.
We find a family of tight Bell inequalities whose coefficients are given by Eq. (6) with $\alpha_{1,\bm{\mu}}=(2,0,-2,1,1,1,1,1,1)$ and $\alpha_{2,\bm{\mu}}=(1,1,1,0,0,0,0,0,0)$, and $n\in\{-2N+1,\,-2N+3,\,\ldots,\,2N-3,\,2N-1\}$ for even $N$ and $n\in\{-2N\,,-2N+2,\,\ldots,\,2N-2,\,2N\}$ for odd $N$, where $\alpha_{1,\bm{\mu}}$ corresponds to the Bell inequality considered in Ref. [47]. Figure 2(c) and (d) show the affine projection and the cross section of this family for $N=11$, respectively. Interestingly, we find that for sufficiently large $N$ the optimal Bell inequality with largest ratio does not belong to this family. For example, for $N=141$, we find $\alpha_{\bm{\mu}}=(694,0,-694,385,309,385,248,309,385)$ with a ratio $\Delta=1.21485$, which exceeds the ratio $\Delta=1.21482$ of $\alpha_{1,\bm{\mu}}$. This implies that with more measurements, finding the Bell inequality with the largest ratio is no longer as straightforward as for $m=2$. Nevertheless, our gradient-based method remains effective in identifying inequalities with larger ratios. In Table LABEL:tab:finiteNwithm=3, we show some other nontrivial tight Bell inequalities not given by Eq. (6).
@@
NN $\Delta_{N,3}$ $\beta_{C}$ $\alpha_{\bm{\mu}}$
@@ 22 $5/4$ 8-8 (2,1,1,2,2,2,2,1,2)(2,1,1,2,2,2,-2,1,-2)
Figure 2: Two-dimensional affine projections (left) and corresponding cross-sections (right) for the $(N,m,2)$ scenarios. (a) and (b) $N=12$, $m=2$. The projection plane is spanned by the two functionals with coefficients $\alpha_{1,\bm{\mu}}=(1,-1,1,1,1)/24$ and $\alpha_{2,\bm{\mu}}=(1,1,0,0,0)/24$. (c) and (d) $N=11$, $m=3$. The affine projection is spanned by $\alpha_{1,\bm{\mu}}=(2,0,-2,1,1,1,1,1,1)/49$ and $\alpha_{2,\bm{\mu}}=(1,1,1,0,0,0,0,0,0)/49$. Dashed lines indicate the maximum ratio in the affine projections and cross sections. Blue dots (left panels) show the vertices in the primal affine projections and correspond to blue lines (right panels) in the dual cross-sections. Cyan lines (left panels) denote tight Bell facets, which map to cyan vertices (right panels).
@@
Conclusion.—We proposed a gradient-based method for optimizing Bell inequalities, demonstrated its effectiveness on PI Bell inequalities involving up to two-body dichotomic correlators, and identified new tight Bell inequalities for the scenario of three measurements per party ($m=3$). The framework is general and can be extended naturally towards certifying quantumness in many-body systems  [18, 19, 20]. It can also be adapted to PI Bell inequalities involving higher-body correlators [31], translationally invariant Bell inequalities [39, 29, 48, 40], and to nonsymmetric Bell inequalities [49, 50, 34, 51, 52]. Here we determine the quantum value as the ground-state energy of its corresponding Hamiltonian within the symmetric measurement setting, serving as a variational upper bound, thereby fixing the Hilbert space dimension. A certified lower bound can be established through SDP-based relaxation methods [53, 54], which we leave for future work.

### v2

@@ 22 $\sqrt{2}$ 2-2 (0,0,1,1,1)(0,0,1,1,-1)
We now apply our method to the $(N,3,2)$ scenario. Since adding a third measurement setting increases the achievable ratio, $\Delta_{N,3}^{\max}\geq\Delta_{N,2}^{\max}$, the corresponding inequalities are more robust and easier to certify experimentally. For $m=3$, all tight Bell inequalities can be enumerated only when $N<5$ [44], and we list these inequalities in [40]. For larger $N$, a complete enumeration is infeasible, but our gradient-based method remains applicable for searching facet Bell inequalities. We find $\Delta_{N,3}^{\max}=\Delta_{N,2}^{\max}$ for small $N$, while for $N\geq 4$ three measurements yield a strictly larger ratio, $\Delta_{N,3}^{\max}>\Delta_{N,2}^{\max}$.
We find a family of tight Bell inequalities whose coefficients are given by Eq. (6) with $\alpha_{1,\boldsymbol{\mu}}=(2,0,-2,1,1,1,1,1,1)$ and $\alpha_{2,\boldsymbol{\mu}}=(1,1,1,0,0,0,0,0,0)$, and $n\in\{-2N+1,\,-2N+3,\,\ldots,\,2N-3,\,2N-1\}$ for even $N$ and $n\in\{-2N\,,-2N+2,\,\ldots,\,2N-2,\,2N\}$ for odd $N$, where $\alpha_{1,\boldsymbol{\mu}}$ corresponds to the Bell inequality considered in Ref. [45]. Figure 2(c) and (d) show the affine projection and the cross section of this family for $N=11$, respectively. Interestingly, we find that for sufficiently large $N$ the optimal Bell inequality with largest ratio does not belong to this family. For example, for $N=141$, we find $\alpha_{\boldsymbol{\mu}}=(694,0,-694,385,309,385,248,309,385)$ with a ratio $\Delta=1.21485$, which exceeds the ratio $\Delta=1.21482$ of $\alpha_{1,\boldsymbol{\mu}}$. This implies that with more measurements, finding the Bell inequality with the largest ratio is no longer as straightforward as for $m=2$. Nevertheless, our gradient-based method remains effective in identifying inequalities with larger ratios (see Table LABEL:tab:finiteNwithm=3).
@@
NN $\Delta_{N,3}$ $\beta_{C}$ $\alpha_{\boldsymbol{\mu}}$
@@ 22 $5/4$ 8-8 (2,1,1,2,2,2,2,1,2)(2,1,1,2,2,2,-2,1,-2)
Figure 2: Two-dimensional affine projections (left) and corresponding cross-sections (right) for the $(N,m,2)$ scenarios. (a) and (b) $N=12$, $m=2$. The projection plane is spanned by the two functionals with coefficients $\alpha_{1,\boldsymbol{\mu}}=(1,-1,1,1,1)/24$ and $\alpha_{2,\boldsymbol{\mu}}=(1,1,0,0,0)/24$. (c) and (d) $N=11$, $m=3$. The affine projection is spanned by $\alpha_{1,\boldsymbol{\mu}}=(2,0,-2,1,1,1,1,1,1)/49$ and $\alpha_{2,\boldsymbol{\mu}}=(1,1,1,0,0,0,0,0,0)/49$. Dashed lines indicate the maximum ratio in the affine projections and cross sections. Blue dots (left panels) show the vertices in the primal affine projections and correspond to blue lines (right panels) in the dual cross-sections. Cyan lines (left panels) denote tight Bell facets, which map to cyan vertices (right panels).
Optimization in the thermodynamic limit $N\rightarrow\infty$.— Permutation symmetry admits a rigorous mean-field solution in the thermodynamic limit. We introduce continuous variables $s_{k}=\langle\mathcal{S}_{k}\rangle/N$ and $z_{kl}=\langle\mathcal{Z}_{kl}\rangle/N$, so that the state concentrates on certain $s_{k}$ and $z_{kl}$ with the two-body correlator approaching $\langle\mathcal{S}_{k}\mathcal{S}_{l}\rangle=N^{2}s_{k}s_{l}$. Equation (5) then becomes $I=N^{2}\sum_{k,l=0}^{m-1}\alpha_{kl}s_{k}s_{l}/2+N\sum_{k=0}^{m-1}\alpha_{k}s_{k}-N\sum_{k,l=0}^{m-1}\alpha_{kl}z_{kl}/2$, where the first term scales as $N^{2}$ and the others as $N$. For the classical case, the local polytope is $\mathcal{N}_{C}=\{(\boldsymbol{s},\boldsymbol{z})|-1\leq s_{k}\leq 1,z_{kk}=1,-1\leq z_{kl}\leq 1,-1+|s_{k}+s_{l}|\leq z_{kl}\leq 1-|s_{k}-s_{l}|\}$, with $\boldsymbol{s}=\{s_{k}\}_{k=0}^{m-1}$ and $\boldsymbol{z}=\{z_{kl}\}_{k,l=0}^{m-1}$. For the quantum case, we find the convex set is given by $\mathcal{N}_{Q}=\{(\boldsymbol{s},\boldsymbol{z})|z_{kk}=1,\;z_{kl}^{2}+s_{k}^{2}+s_{l}^{2}-2z_{kl}s_{k}s_{l}\leq 1\}$. The classical bound and the minimal quantum value in the thermodynamic limit can be written as the optimization over the two sets.
The ratios in the $N\rightarrow\infty$ limit are summarized in Table LABEL:tab:infiniteN-different-m. We consider two optimizations: (i) fixing all $\alpha_{kl}=1$ and optimizing only the single-body coefficients $\alpha_{k}$; (ii) optimizing both $\alpha_{kl}$ and $\alpha_{k}$, which yields larger ratios for finite $m$. The large $m$ admits the same results $\Delta_{\infty,\infty}=\coth(1)\approx 1.31304$. The detailed derivations is left in [40].
Table 3: Ratios for the scenario $(N,m,2)$ in the limit $N\rightarrow\infty$. The second column shows $\Delta_{\infty,m}$ with fixed $\alpha_{kl}=1$ (only $\alpha_{k}$ optimized), and the third column shows the maximum ratio $\Delta_{\infty,m}^{\mathrm{max}}$ when both $\alpha_{kl}$ and $\alpha_{k}$ are optimized.
$m$ $\Delta_{\infty,m}$ $\Delta_{\infty,m}^{\mathrm{max}}$
2 $5/4$ $5/4$
3 $1.28458$ $9/7$
4 $1.29711$ $353/272$
5 $1.30285$ $275/211$
6 $1.30597$ $66637/51012$
$\infty$ $\coth(1)$ $\coth(1)$
Conclusion.—We proposed a gradient-based method for optimizing Bell inequalities, applied it to PI Bell inequalities involving up to two-body dichotomic correlators, and identified new tight Bell inequalities for the scenario of three measurements per party ($m=3$). We also derive analytical expressions for the maximum ratio in the thermodynamic limit $N\to\infty$, which yield simple integer ratios for small $m$ and converge to $\coth(1)$ in the joint limit $N,m\to\infty$.
@@ 2 $5/4$ $5/4$
The framework is general and naturally extends to certifying quantumness in many-body systems  [18, 19, 20, 46]. In particular, one can fix a Bell operator and inversely optimize the classical bound, leading to a more robust certification of nonlocality for the given quantum state [20, 47]. It can also be adapted to PI Bell inequalities involving higher-body correlators [30], translationally invariant Bell inequalities [48, 28, 49, 50, 51, 52], and other more general Bell inequalities [53, 54, 33, 55, 56]. Here we determine the quantum value as the ground-state energy of its corresponding Hamiltonian within the symmetric measurement setting, providing a variational upper bound at fixed Hilbert-space dimension. A certified lower bound can be established through SDP-based relaxation methods [57, 58, 59], which we leave for future work.